OpenStudy (anonymous):

A toy rocket was launched from the ground. The function f(x) = -16x2 + 128x shows the height of the rocket f(x), in feet, from the ground at time x seconds. What is the axis of symmetry of the graph of f(x), and what does it represent?

OpenStudy (anonymous):

so the 'x' coordinate of our vertex is 4... so I'm going to say the rocket would take 4 seconds to get up to maximum height...and then it will fall back down taking 4 seconds How do i know it will only take 4 seconds to get down? Well I would find out the x-intercepts of the parabola by factoring -16x^2 + 128x = 0 factor our like terms -16x(x - 8) = 0 so we have 2 equations -16x = 0 and (x - 8) = 0 well that means our 2 x intercepts are 0 and 8 So if it take 4 seconds to go up to the max height...we only have 4 seconds to fall back down Hence the answer being B

OpenStudy (anonymous):

Woah, Where did you get that answer from?

OpenStudy (misty1212):

what on earth?

OpenStudy (misty1212):

first coordinate of the vertex is always (i repeat always ) \[-\frac{b}{2a}\]

OpenStudy (misty1212):

do \[-\frac{128}{2\times (-16)}\] and that is all

OpenStudy (anonymous):

@misty1212 What he did was, He went on google, Searched up my question. And some other guy that has already posted this question answered it.

OpenStudy (misty1212):

i love google, but \(-\frac{b}{2a}\) cannot be that hard to compute in any case what does "the answer is B" mean in this context?

OpenStudy (anonymous):

Like this http://widget.openstudy.com/study#/updates/5344393ae4b02bb88e656693 Then he just takes credit for the answer

OpenStudy (anonymous):

Thanks @LesTwins I do the same. I already got that answer but I want to double check. Quite slick.

OpenStudy (anonymous):

I did the math then searched with google to give u the insurance you need to get the question correct